# SESR-Eval: Dataset for Evaluating LLMs in the Title-Abstract Screening of Systematic Reviews

### v1

@@ IV-B What is the performance of the LLMs in the screening task?
Llama 4 Maverick Primary (n=34,528) 0.87 0.47 0.61 0.58
Secondary (n=24) 0.74 0.54 0.61 0.50
@@
Secondary (n=24) 0.73 0.60 0.49 0.46
@@
Secondary (n=24) 0.73 0.53 0.66 0.51
@@
Secondary (n=24) 0.73 0.53 0.60 0.48
@@
Secondary (n=24) 0.73 0.62 0.38 0.41
@@
Secondary (n=24) 0.59 0.39 0.76 0.45
@@
Secondary (n=24) 0.73 0.58 0.45 0.44
@@
Secondary (n=24) 0.72 0.58 0.41 0.41
@@
Secondary (n=24) 0.34 0.34 1.00 0.44
@@
Accuracy Primary GPT-4.1 mini
Accuracy Secondary
Llama4 Maverick
o3-mini
GPT-4.1 mini
Claude 3.7 Sonnet
F1 Primary Llama 4 Maverick
F1 Secondary Llama4 Maverick, GPT-4o
Table VII also reports precision and recall for the binary decisions made by the models. To highlight differences in precision and recall, we present Likert scale results in Fig. 3, where the precision, recall, and F1 score curves at each decision point provide a clearer picture. The Likert scale is described in Section III-C. From Fig. 3 we observe that the F1 score behaves as expected, forming an inverted U-shape and peaking in the middle of the Likert scale. Recall decreases as we move along the X-axis from left to right, dropping from a perfect 1.001.001.00 to 0.230.230.23. Precision exhibits the opposite trend, increasing from 0.120.120.12 to 0.500.500.50.
Often, a researcher using the LLMs wants to optimize for maximum recall, as missing evidence (false negatives) is more difficult to recover from than incorrectly including a paper (false positive). Here, we assume a process where an LLM performs the initial title-abstract screening, and then the included papers are checked by the researcher. In such a setup, a deemed useful LLM could be defined as follows. We would aim for high recall, at least above 0.950.950.95, with reasonable precision, around 0.500.500.50. This would mean capturing 95% of the evidence, while manually screening papers of which 50% contain relevant evidence. Unfortunately, as seen in Fig. 3, such a point is not reached with current LLMs.
11122233344455566677700.20.20.20.40.40.40.60.60.60.80.80.8111Likert-scalePrecisionRecallF1
@@ IV-C How does screening performance vary across secondary studies?
Table IX shows the accuracy, precision, recall, and F1 score for the secondary studies across all LLMs. We observe notable variation between the secondary studies, with accuracy ranging from 0.340.340.34 to 0.850.850.85 and F1 scores ranging from 0.070.070.07 to 0.920.920.92. In no study do we reach our deemed useful thresholds for LLMs, as none achieve a recall above 0.95 while maintaining a precision of 0.500.500.50.
i [42]
11122233344455566677700.20.20.20.40.40.40.60.60.60.80.80.8111PrecisionRecallF1
ii [44]
11122233344455566677700.20.20.20.40.40.40.60.60.60.80.80.8111PrecisionRecallF1
iii [53]
11122233344455566677700.20.20.20.40.40.40.60.60.60.80.80.8111PrecisionRecallF1
iv [57]
11122233344455566677700.20.20.20.40.40.40.60.60.60.80.80.8111PrecisionRecallF1
v [62]
11122233344455566677700.20.20.20.40.40.40.60.60.60.80.80.8111PrecisionRecallF1
@@
Fig. 4 shows the precision and recall curves for secondary studies that have a vastly different curve shape than in Fig. 3. We notice that for the study [53], there is actually a point on the curve where the deemed useful threshold is reached. However, it is important to note that this is likely because the dataset for that paper contains only included studies, resulting in perfect precision and recall when all papers are included at the lowest Likert scores. If we consider this study an outlier, we are left with no cases where the deemed useful thresholds are met.
Looking at the F1 curves also reveals variability between studies. Most studies have the highest F1 scores in the middle of the Likert scale, as one would expect. However, there are multiple studies where the F1 score drops when moving from left to right along the X-axis [42, 44, 53, 57, 62]. These studies would clearly benefit from using the lowest values on the Likert scale and suffer from the binary decision threshold used in our experiment, which mostly corresponds to the median value on the Likert scale—number 4 in the figure.
@@
[41] 0.78 0.80 0.86 0.83
[42] 0.34 0.83 0.28 0.42
[43] 0.66 0.49 0.55 0.51
[44] 0.66 0.66 0.73 0.69
[45] 0.72 0.13 0.76 0.22
[46] 0.80 0.03 0.57 0.07
[47] 0.78 0.28 0.75 0.41
[48] 0.63 0.39 0.80 0.52
[49] 0.79 0.10 0.78 0.18
[50] 0.67 0.14 0.82 0.23
[51] 0.76 0.39 0.55 0.45
[52] 0.83 0.08 0.40 0.14
[53] 0.85 1.00 0.85 0.92
[54] 0.80 0.16 0.35 0.22
[55] 0.64 0.65 0.50 0.57
[56] 0.59 0.56 0.82 0.66
[57] 0.44 0.76 0.32 0.45
[58] 0.59 0.42 0.57 0.48
[59] 0.58 0.14 0.36 0.20
[60] 0.82 0.13 0.46 0.21
[61] 0.71 0.22 0.64 0.32
[62] 0.56 0.72 0.42 0.53
[63] 0.75 0.37 0.76 0.50
[64] 0.40 1.00 0.40 0.57
@@ IV-D What is the combined effect of LLMs and secondary studies on the screening results?
Reference - [53]
[41] -0.702 0.145 -4.85 < 2e-16 *** 0.50
[42] -3.240 0.151 -21.50 < 2e-16 *** 0.04
[43] -1.599 0.148 -10.82 < 2e-16 *** 0.20
[44] -1.621 0.149 -10.87 < 2e-16 *** 0.20
[45] -1.194 0.142 -8.39 < 2e-16 *** 0.30
[46] -0.493 0.143 -3.45 0.000570 *** 0.61
[47] -0.697 0.144 -4.84 1.33e-06 *** 0.50
[48] -1.775 0.146 -12.16 < 2e-16 *** 0.17
[49] -0.665 0.143 -4.63 3.66e-06 *** 0.51
[50] -1.554 0.149 -10.40 < 2e-16 *** 0.21
[51] -0.920 0.145 -6.29 3.12e-10 *** 0.40
[52] -0.183 0.144 -1.27 0.20 0.83
[54] -0.510 0.143 -3.56 0.000376 *** 0.60
[55] -1.720 0.149 -11.55 < 2e-16 *** 0.18
[56] -2.017 0.160 -12.63 < 2e-16 *** 0.13
[57] -2.760 0.146 -18.91 < 2e-16 *** 0.06
[58] -1.989 0.154 -12.91 < 2e-16 *** 0.14
[59] -2.024 0.149 -13.60 < 2e-16 *** 0.13
[60] -0.271 0.143 -1.89 0.06 0.76
[61] -1.264 0.144 -8.79 < 2e-16 *** 0.28
[62] -2.144 0.148 -14.46 < 2e-16 *** 0.12
[63] -0.931 0.146 -6.38 1.81e-10 *** 0.39
[64] -2.942 0.159 -18.48 < 2e-16 *** 0.05
@@
As both the LLMs and secondary studies impact the results of LLM-based paper screening, we decided to investigate how these factors work together. We did this by fitting a logistic regression model predicting the correctness of screening decisions based on the LLM used and the secondary study. We selected the best-performing model and secondary study for reference categories, as this makes the resulting model easier to interpret. The model is shown in Table X. In the model, a statistically significant intercept (log-odds = 3.203.203.20, p < $2\mathrm{e}{-16}$) corresponds to high baseline odds of a correct prediction when using the reference model (GPT-4.1 mini) and reference study ([53]) (odds ratio 24.08$\approx$ 24.08≈ 24.08).
@@
As the best performing LLM was selected to be the reference category, all other LLMs have reduced odds of correctness relative to the baseline. Both Ministral 8B and GPT-4.1 nano showed a notably large negative effect ($\beta$ = 4.332-4.332- 4.332, p < $2\mathrm{e}{-16}$, OR 0.013$\approx$ 0.013≈ 0.013) ($\beta$ = 2.05-2.05- 2.05, p < $2\mathrm{e}{-16}$, OR 0.128$\approx$ 0.128≈ 0.128), suggesting substantially poorer performance compared to the larger LLMs that had notable better coefficient and odds-rations.
@@
Other LLMs such as GPT-4o (OR 0.58$\approx$ 0.58≈ 0.58) and GPT-4.1 (OR 0.54$\approx$ 0.54≈ 0.54) also demonstrated significantly diminished performance. However, LLMs such as Claude 3.7 Sonnet and DeepSeek R1 were close to the performance of the reference LLM (ORs 0.930.94$\approx$ 0.93-0.94≈ 0.93 - 0.94).
@@
Regarding the influence of secondary studies, the regression model identified considerable variability in the odds of correct classification across sources. Several secondary studies were associated with large negative effects. For example, screening tasks from Alonso et al. [42] and Tebes et al.[57] exhibited strong reductions in predictive accuracy (ORs 0.05$\approx$ 0.05≈ 0.05 and 0.060.060.06, respectively). Conversely, a few studies (e.g., Rani et al. [52] and Kuutila et al.[60], OR 0.83$\approx$ 0.83≈ 0.83 and 0.780.780.78) had relatively weaker effects. Thus, both the LLM used and the characteristics of the secondary study significantly impact the likelihood of correct screening in our dataset, when using zero-shot prompts.
@@ IV-E How does the time, token efficiency, and cost vary across different LLMs?
In terms of time efficiency, we found that GPT-4.1 mini and GPT-4o achieved performance comparable with slower models (see Table VII), while offering a three- to five-fold reduction in screening time per paper compared to the highest-performing models (see Table XI). Among the three fastest models, Ministral 8B and GPT-4.1 nano – screening performance was notably limited. For comparison, Huotala et al. [3] reported that expert human screeners required an average of 85.95 seconds to screen a single paper. In contrast, every LLM except o3-mini were faster than the human screener.
@@
[…]
@@
LLM
Input
\$ / 1M
Tokens
Output
\$ / 1M
Tokens
Input tokens (per study) Output tokens (per study) Time (P95) Per paper
Cost per
secondary
study

### v2

@@ IV-B What is the performance of the LLMs in the screening task?
Llama 4 Maverick Primary (n=34,528) 0.87 0.47 0.61 0.53
Secondary (n=24) 0.74 0.54 0.61 0.50111Average F1 score of the F1 scores from the 24 secondary studies. 0.57222F1 score calculated from the average precision and recall.
@@
Secondary (n=24) 0.73 0.60 0.49 0.46111Average F1 score of the F1 scores from the 24 secondary studies. 0.54222F1 score calculated from the average precision and recall.
@@
Secondary (n=24) 0.73 0.53 0.66 0.51111Average F1 score of the F1 scores from the 24 secondary studies. 0.59222F1 score calculated from the average precision and recall.
@@
Secondary (n=24) 0.73 0.53 0.60 0.48111Average F1 score of the F1 scores from the 24 secondary studies. 0.56222F1 score calculated from the average precision and recall.
@@
Secondary (n=24) 0.73 0.62 0.38 0.41111Average F1 score of the F1 scores from the 24 secondary studies. 0.47222F1 score calculated from the average precision and recall.
@@
Secondary (n=24) 0.59 0.39 0.76 0.45111Average F1 score of the F1 scores from the 24 secondary studies. 0.52222F1 score calculated from the average precision and recall.
@@
Secondary (n=24) 0.73 0.58 0.45 0.44111Average F1 score of the F1 scores from the 24 secondary studies. 0.51222F1 score calculated from the average precision and recall.
@@
Secondary (n=24) 0.72 0.58 0.41 0.41111Average F1 score of the F1 scores from the 24 secondary studies. 0.48222F1 score calculated from the average precision and recall.
@@
Secondary (n=24) 0.34 0.34 1.00 0.44111Average F1 score of the F1 scores from the 24 secondary studies. 0.51222F1 score calculated from the average precision and recall.
@@
Accuracy Primary study GPT-4.1 mini
Accuracy Secondary study Llama 4 Maverick
F1 Primary study Llama 4 Maverick
F [^1] footnotemark: 1, F122footnotemark: 2 Secondary study GPT-4o
Table VII also reports precision and recall for the binary decisions made by the models. To highlight differences in precision and recall, we present Likert scale results in Fig. 3, where the precision, recall, and F1 score curves at each decision point provide a clearer picture. The Likert scale is described in Section III-C. From Fig. 3 we observe that the F1 score behaves as expected, forming an inverted U-shape and peaking in the middle of the Likert scale. Recall decreases as we move along the X-axis from left to right, dropping from a perfect $1.00$ to $0.23$. Precision exhibits the opposite trend, increasing from $0.12$ to $0.50$.
Often, a researcher using the LLMs wants to optimize for maximum recall, as missing evidence (false negatives) is more difficult to recover from than incorrectly including a paper (false positive). Here, we assume a process where an LLM performs the initial title-abstract screening, and then the included papers are checked by the researcher. In such a setup, a deemed useful LLM could be defined as follows. We would aim for high recall, at least above $0.95$, with reasonable precision, around $0.50$. This would mean capturing 95% of the evidence, while manually screening papers of which 50% contain relevant evidence. Unfortunately, as seen in Fig. 3, such a point is not reached with current LLMs.
1122334455667700.20.20.40.40.60.60.80.811Likert-scalePrecisionRecallF1
@@ IV-C How does screening performance vary across secondary studies?
Table IX shows the accuracy, precision, recall, and F1 score for the secondary studies across all LLMs. We observe notable variation between the secondary studies, with accuracy ranging from $0.34$ to $0.85$ and F1 scores ranging from $0.07$ to $0.92$. In no study do we reach our deemed useful thresholds for LLMs, as none achieve a recall above 0.95 while maintaining a precision of $0.50$.
i [alonsoSystematicMappingStudy2023b]
1122334455667700.20.20.40.40.60.60.80.811PrecisionRecallF1
$i$ [sharbafConflictManagementTechniques2023a]
1122334455667700.20.20.40.40.60.60.80.811PrecisionRecallF1
iii [pereiraLearningSoftwareConfiguration2021a]
1122334455667700.20.20.40.40.60.60.80.811PrecisionRecallF1
iv [tebesAnalyzingDocumentingSystematic2020a]
1122334455667700.20.20.40.40.60.60.80.811PrecisionRecallF1
v [teixeiraModelingAutomaticCode2017a]
1122334455667700.20.20.40.40.60.60.80.811PrecisionRecallF1
@@
Fig. 4 shows the precision and recall curves for secondary studies that have a vastly different curve shape than in Fig. 3. We notice that for the study [pereiraLearningSoftwareConfiguration2021a], there is actually a point on the curve where the deemed useful threshold is reached. However, it is important to note that this is likely because the dataset for that paper contains only included studies, resulting in perfect precision and recall when all papers are included at the lowest Likert scores. If we consider this study an outlier, we are left with no cases where the deemed useful thresholds are met.
Looking at the F1 curves also reveals variability between studies. Most studies have the highest F1 scores in the middle of the Likert scale, as one would expect. However, there are multiple studies where the F1 score drops when moving from left to right along the X-axis [alonsoSystematicMappingStudy2023b, sharbafConflictManagementTechniques2023a, pereiraLearningSoftwareConfiguration2021a, tebesAnalyzingDocumentingSystematic2020a, teixeiraModelingAutomaticCode2017a]. These studies would clearly benefit from using the lowest values on the Likert scale and suffer from the binary decision threshold used in our experiment, which mostly corresponds to the median value on the Likert scale—number 4 in the figure.
@@
[stradowskiMachineLearningSoftware2023a] 0.78 0.80 0.86 0.83
[alonsoSystematicMappingStudy2023b] 0.34 0.83 0.28 0.42
[somersDigitaltwinbasedTestingCyber2023b] 0.66 0.49 0.55 0.51
[sharbafConflictManagementTechniques2023a] 0.66 0.66 0.73 0.69
[zakeri-nasrabadiSystematicLiteratureReview2023b] 0.72 0.13 0.76 0.22
[bjarnasonEmpiricallyBasedModel2023a] 0.80 0.03 0.57 0.07
[lewowskiHowFarAre2022a] 0.78 0.28 0.75 0.41
[vandinterPredictiveMaintenanceUsing2022a] 0.63 0.39 0.80 0.52
[casalaroModeldrivenEngineeringMobile2022a] 0.79 0.10 0.78 0.18
[deckersSystematicLiteratureReview2022a] 0.67 0.14 0.82 0.23
[tambonHowCertifyMachine2022a] 0.76 0.39 0.55 0.45
[raniDecadeCodeComment2023a] 0.83 0.08 0.40 0.14
[pereiraLearningSoftwareConfiguration2021a] 0.85 1.00 0.85 0.92
[rodriguez-perezPerceivedDiversitySoftware2021a] 0.80 0.16 0.35 0.22
[barisicMultiparadigmModelingCyber2022a] 0.64 0.65 0.50 0.57
[ebrahimiMobileAppPrivacy2021a] 0.59 0.56 0.82 0.66
[tebesAnalyzingDocumentingSystematic2020a] 0.44 0.76 0.32 0.45
[geraldiSoftwareProductLine2020a] 0.59 0.42 0.57 0.48
[lenarduzziSystematicLiteratureReview2021a] 0.58 0.14 0.36 0.20
[kuutilaTimePressureSoftware2020b] 0.82 0.13 0.46 0.21
[shevtsovControlTheoreticalSoftwareAdaptation2018a] 0.71 0.22 0.64 0.32
[teixeiraModelingAutomaticCode2017a] 0.56 0.72 0.42 0.53
[linOpinionMiningSoftware2022a] 0.75 0.37 0.76 0.50
[zhangTestingVerificationNeuralnetworkbased2020a] 0.40 1.00 0.40 0.57
@@ IV-D What is the combined effect of LLMs and secondary studies on the screening results?
Reference - [pereiraLearningSoftwareConfiguration2021a]
[stradowskiMachineLearningSoftware2023a] -0.702 0.145 -4.85 < 2e-16 *** 0.50
[alonsoSystematicMappingStudy2023b] -3.240 0.151 -21.50 < 2e-16 *** 0.04
[somersDigitaltwinbasedTestingCyber2023b] -1.599 0.148 -10.82 < 2e-16 *** 0.20
[sharbafConflictManagementTechniques2023a] -1.621 0.149 -10.87 < 2e-16 *** 0.20
[zakeri-nasrabadiSystematicLiteratureReview2023b] -1.194 0.142 -8.39 < 2e-16 *** 0.30
[bjarnasonEmpiricallyBasedModel2023a] -0.493 0.143 -3.45 0.000570 *** 0.61
[lewowskiHowFarAre2022a] -0.697 0.144 -4.84 1.33e-06 *** 0.50
[vandinterPredictiveMaintenanceUsing2022a] -1.775 0.146 -12.16 < 2e-16 *** 0.17
[casalaroModeldrivenEngineeringMobile2022a] -0.665 0.143 -4.63 3.66e-06 *** 0.51
[deckersSystematicLiteratureReview2022a] -1.554 0.149 -10.40 < 2e-16 *** 0.21
[tambonHowCertifyMachine2022a] -0.920 0.145 -6.29 3.12e-10 *** 0.40
[raniDecadeCodeComment2023a] -0.183 0.144 -1.27 0.20 0.83
[rodriguez-perezPerceivedDiversitySoftware2021a] -0.510 0.143 -3.56 0.000376 *** 0.60
[barisicMultiparadigmModelingCyber2022a] -1.720 0.149 -11.55 < 2e-16 *** 0.18
[ebrahimiMobileAppPrivacy2021a] -2.017 0.160 -12.63 < 2e-16 *** 0.13
[tebesAnalyzingDocumentingSystematic2020a] -2.760 0.146 -18.91 < 2e-16 *** 0.06
[geraldiSoftwareProductLine2020a] -1.989 0.154 -12.91 < 2e-16 *** 0.14
[lenarduzziSystematicLiteratureReview2021a] -2.024 0.149 -13.60 < 2e-16 *** 0.13
[kuutilaTimePressureSoftware2020b] -0.271 0.143 -1.89 0.06 0.76
[shevtsovControlTheoreticalSoftwareAdaptation2018a] -1.264 0.144 -8.79 < 2e-16 *** 0.28
[teixeiraModelingAutomaticCode2017a] -2.144 0.148 -14.46 < 2e-16 *** 0.12
[linOpinionMiningSoftware2022a] -0.931 0.146 -6.38 1.81e-10 *** 0.39
[zhangTestingVerificationNeuralnetworkbased2020a] -2.942 0.159 -18.48 < 2e-16 *** 0.05
@@
As both the LLMs and secondary studies impact the results of LLM-based paper screening, we decided to investigate how these factors work together. We did this by fitting a logistic regression model predicting the correctness of screening decisions based on the LLM used and the secondary study. We selected the best-performing model and secondary study for reference categories, as this makes the resulting model easier to interpret. The model is shown in Table X. In the model, a statistically significant intercept (log-odds = $3.20$, p < $2\mathrm{e}{-16}$) corresponds to high baseline odds of a correct prediction when using the reference model (GPT-4.1 mini) and reference study ([pereiraLearningSoftwareConfiguration2021a]) (odds ratio $\approx 24.08$).
@@
As the best performing LLM was selected to be the reference category, all other LLMs have reduced odds of correctness relative to the baseline. Both Ministral 8B and GPT-4.1 nano showed a notably large negative effect ($\beta$ = 4.332-4.332, p < $2\mathrm{e}{-16}$, OR $\approx 0.013$) ($\beta$ = 2.05-2.05, p < $2\mathrm{e}{-16}$, OR $\approx 0.128$), suggesting substantially poorer performance compared to the larger LLMs that had notable better coefficient and odds-rations.
@@
Other LLMs such as GPT-4o (OR $\approx 0.58$) and GPT-4.1 (OR $\approx 0.54$) also demonstrated significantly diminished performance. However, LLMs such as Claude 3.7 Sonnet and DeepSeek R1 were close to the performance of the reference LLM (ORs $\approx 0.93-0.94$).
@@
Regarding the influence of secondary studies, the regression model identified considerable variability in the odds of correct classification across sources. Several secondary studies were associated with large negative effects. For example, screening tasks from Alonso et al. [alonsoSystematicMappingStudy2023b] and Tebes et al.[tebesAnalyzingDocumentingSystematic2020a] exhibited strong reductions in predictive accuracy (ORs $\approx 0.05$ and $0.06$, respectively). Conversely, a few studies (e.g., Rani et al. [raniDecadeCodeComment2023a] and Kuutila et al.[kuutilaTimePressureSoftware2020b], OR $\approx 0.83$ and $0.78$) had relatively weaker effects. Thus, both the LLM used and the characteristics of the secondary study significantly impact the likelihood of correct screening in our dataset, when using zero-shot prompts.
@@ IV-E How does the time, token efficiency, and cost vary across different LLMs?
In terms of time efficiency, we found that GPT-4.1 mini and GPT-4o achieved performance comparable with slower models (see Table VII), while offering a three- to five-fold reduction in screening time per paper compared to the highest-performing models (see Table XI). Among the three fastest models, Ministral 8B and GPT-4.1 nano – screening performance was notably limited. For comparison, Huotala et al. [huotalaPromiseChallengesUsing2024a] reported that expert human screeners required an average of 85.95 seconds to screen a single paper. In contrast, every LLM except o3-mini were faster than the human screener.
@@
Regarding costs, we observe that even the most expensive model (o3-mini) cost only $24*\mathdollar 36.7\approx\mathdollar 881$ across the 24 secondary studies. As such, it appears that the cost of using LLMs should not be a barrier to their use in title-abstract screening. When it comes to cost-efficiency, we find that Meta’s LLaMA 4 Maverick (cost per secondary study \$1.4) and GPT-4.1 mini (cost per secondary study \$2.2) delivered performance on par with more expensive models (see Table VII), while offering over tenfold savings in cost compared to the more expensive models (see Table XI).
@@
LLM Input \$ / 1M Tokens Output \$ / 1M Tokens Input tokens (per study) Output tokens (per study) Time (P95) Per paper Cost per secondary study